# Gravitational Lensing of Euler-Heisenberg Black Hole Surrounded by Perfect Fluid Dark Matter

### v1

@@ Appendix A Horizon of Euler-Heisenberg Black Hole Surrounded by Perfect Fluid Dark Matter
When taking into account both nonlinear electrodynamics effects (with $a>0$) and dark matter effects (with $\lambda_{\text{DM}}\neq 0$), the horizons of Euler-Heisenberg black hole surrounded by PFDM are plotted in the lower panels of figure 8, for nonlinear electrodynamics parameters $a=0.2M^{2}$ and $a=0.8M^{2}$ respectively. Comparing the positions of horizon for Euler-Heisenberg black hole without dark matter (in the upper right panel) and those in the presence of dark matter (in the lower left and lower right panels), we can obtain the following conclusions. In the presence of dark matter and nonlinear electrodynamics, the naked singularity is no longer emerged for any black hole electric charge, which is the same as that for Euler-Heisenberg black hole without PFDM. The dark matter influences the horizon size and changes the wide of three horizon region in the parameter space of black hole charge. The negative PFDM parameter ($\lambda_{\text{DM}}<0$) results in large horizon radius, which can be explained using the “effective mass” role the PFDM played in the spacetime metric. For an intermediate electric charge, a positive PFDM parameter ($\lambda_{\text{DM}}>0$) narrows the three horizon region, while a negative PFDM parameter ($\lambda_{\text{DM}}<0$) widens the three horizon region. For instance, when nonlinear electrodynamics parameter gets value $a=0.2M^{2}$, the PFDM with $\lambda_{\text{DM}}=-0.5M$ produces a much broader three horizon region, compared with those without PFDM plotted in the upper right panel. When $a=0.8M^{2}$, the Euler-Heisenberg black hole in the absence of dark matter does not possess a three horizons region, while the three horizon region emerges for negative PFDM parameter values $\lambda_{\text{DM}}=-0.2M$ and $\lambda_{\text{DM}}=-0.5M$.

### v2

@@ Appendix A Horizons for Euler-Heisenberg Black Hole Surrounded by Perfect Fluid Dark Matter
When taking into account both nonlinear electrodynamics effects (with $a>0$) and dark matter effects (with $\lambda_{\text{DM}}>0$), the horizons of Euler-Heisenberg black hole surrounded by PFDM are plotted in the lower panels of figure 8, for nonlinear electrodynamics parameters $a=0.2M^{2}$ and $a=0.8M^{2}$ respectively. Comparing the positions of horizon for Euler-Heisenberg black hole without dark matter (in the upper right panel) and those in the presence of dark matter (in the lower left and lower right panels), we can obtain the following conclusions. In the presence of dark matter and nonlinear electrodynamics, the naked singularity is no longer emerged for arbitrary black hole electric charge, which is the same as that for Euler-Heisenberg black hole without PFDM. The dark matter influences the horizon size and changes the wide of three horizon region in the parameter space of black hole charge. A positive PFDM parameter ($\lambda_{\text{DM}}>0$) results in smaller horizon radius, which can be explained using the “effective mass” role the PFDM played in spacetime metric. Furthermore, a positive PFDM parameter narrows the three horizon region in the parameter space of black hole charge. For instance, when nonlinear electrodynamics parameter gets value $a=0.2M^{2}$, the three horizon region get narrowed for $0<\lambda_{\text{DM}}\leq 0.3M$ (compared with those without PFDM plotted in the upper right panel), while the $\lambda_{\text{DM}}=0.5M$ case does not possess a three horizon region. For a large nonlinear electrodynamics parameter (such as $a=0.8M^{2}$ in the lower right panel), the Euler-Heisenberg black hole in the presence of dark matter does not possess a three horizons region, no matter what the PFDM parameter is varied.